OpenStudy (help!!!!):

Solve x+ square root x = 5

OpenStudy (anonymous):

\[x+\sqrt{x}\]=5 first of all, lets square everything to get rid of the square root so it would be \[x^2 + x = 5^2\] now lets move everything to one side : \[x^2 + x - 25 = 0\] so now, it can be solved in the normal way of solving quadratic equations